# Continuous Audio Language Models

### v2

@@ 3. Consistency-Model Head
$\small\mathcal{L}_{\text{CALM}}(\theta,\phi,\psi)=\sum_{s=1}^{S}\mathbb{E}_{t,\bm{\epsilon}}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-F_{\bar{\phi}}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-\cos(t)\frac{df_{\bar{\phi}}(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s})}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right],$ (3)
@@ 4.3 Combining noisy long-term context and clean short-term context
In preliminary experiments, we realized that the music generation model was diverging quite quickly during inference because it was not robust to error accumulation. Pasini et al. (2024b) introduce a noise augmentation trick at training time in order to tackle this problem. Given a sequence $(\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{S})$, they sample $k_{s}\sim\mathcal{U}(0,1)$ and $\bm{\epsilon}_{s}\sim\mathcal{N}(0,\mathbf{I})$ for every $s\in\{1,\ldots,S\}$ and compute $\tilde{\mathbf{x}}^{s}=k_{s}\bm{\epsilon}_{s}+(1-k_{s})\mathbf{x}^{s}$ for every $s$. After comparison and to preserve the variance of the $\mathbf{x}^{s}$ that are gaussians, we modify the noising formula as follows: $\tilde{\mathbf{x}}^{s}=\sqrt{k_{s}}\bm{\epsilon}_{s}+\sqrt{1-k_{s}}\mathbf{x}^{s}$
In our case, applying this noising method to music generation slightly improves model stability but often reduces detail and instrument diversity, typically preserving only rhythmic elements, with audio fading into silence after 10–15 seconds. Since Pasini et al. (2024b) targets short, single-instrument clips, it’s unsurprising the method performs best in that constrained setting. We hypothesize that the added noise inhibits the backbone transformer from encoding fine-grained information into $\mathbf{z}^{s}_{\text{long}}$, limiting the MLP’s ability to reconstruct detailed audio. However, combining this with a short-context transformer computing $\mathbf{z}^{s}_{\text{short}}$ yields the best results (Tab. 5), likely because the clean short-term context restores local detail needed to model the distribution of the next $\mathbf{x}^{s}$.
At inference time, we omit noise augmentation and achieve our best results with this setup. Providing historical context to the diffusion head was used in the TTS model DiTAR (Jia et al., 2025), which reported a significant reduction in word error rate when using it.
@@ 4.5 Gaussian temperature sampling
Sampling strategies, such as temperature sampling, have a significant impact on generation quality in the discrete setup, particularly for speech. To replicate this behavior in the continuous domain, we introduce a sampling heuristic that results in similar gains.
Let $\bm{\epsilon}\sim\mathcal{N}(0,I)$ be a standard multivariate Gaussian random variable and let $\mathbf{x}=f(\bm{\epsilon})$, where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a differentiable function. By the change of variables formula, the probability density of $\mathbf{x}$ is given by: $p(\mathbf{x})=p(\bm{\epsilon})\cdot\left|\det\left(\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}\right)\right|^{-1}$, where $\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}$ is the Jacobian matrix of $f$ at $\bm{\epsilon}$. In theory, to apply temperature sampling with temperature $\tau$ we should raise both terms to the power $1/\tau$, but the Jacobian term is untractable. In practice, we noticed that we can apply temperature only on the Gaussian noise sampling (which is equivalent to sampling from a Gaussian with variance $1/\sqrt{\tau}$) and get significant improvements in metrics.
The effects of gaussian temperature are further discussed in Section 5.3.
Table 1: Speech compression models. Our VAE is on par with the VQ-VAE on acoustic quality (MOSNeT) and outperforms it on semantic discriminability (ABX).
Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbit/s) MOSNet ($\uparrow$) ABX ($\downarrow$)
VQ-VAE (Mimi) 8 RVQ 12.5 1.1kbps 3.11 9.4%
VAE 32 dims 12.5 - 3.15 8.1%
@@ 5 Experiments and results
We now present experimental results on both speech and music continuation. We generate 27 seconds following a prompt of 3 seconds. All detailed hyperparameters are in the Appendix Sec. C. All metrics reported with confidence intervals use 95% confidence. Metrics without confidence intervals are metrics computed over the entire test dataset.
@@ 5.1 Speech Continuation
Results: Tab. 1 shows that our 32-dimensional VAE matches an 8-RVQ Mimi codec on MOSNet (Lo et al., 2019), which measures audio quality, and exceeds it on the ABX metric (Schatz et al., 2013). ABX evaluates phonetic discriminability by testing whether a word like “bat” is represented closer to another “bat” utterance than to a similar-sounding word like “bit”, based on latent distances.
Tab. 2 shows that the 1-step Consistency model outperforms the RQ-Transformer with 8 RVQ on all our automatic and human based metrics as well as on speed. For automatic metrics, we compute PPX and VERT Lakhotia et al. (2021). The PPX metric measures the semantic meaningfulness of the generated speech. To do so we generate 1000 excerpts of speech of 30 second, we use Whisper Radford et al. (2022) to compute textual transcriptions and finally compute the negative log-likelihood of the text tokens with a Mistral 7B LLM Jiang et al. (2023) and convert it to Perplexity. Because a model that generates poorly diverse but good quality sentences would perform well on the PPX metric, the authors of Lakhotia et al. (2021) introduce the VERT metric (for diVERsiTy) which is a geometric mean of self- and auto-BLEU metrics. We use the official implementation from the fairseq Ott et al. (2019) repository.
To assess perceptual quality, we conduct two human evaluation studies involving 50 participants and 50 randomly selected examples from the English test set. Each participant rates 10 examples across the following evaluation protocols: For Acoustic Quality, participants are presented with all model continuations for the same prompt, including the ground truth reference, and rate the acoustic quality of each continuation on a 1 to 5 scale. For Meaningfulness, participants are shown two continuations of the same prompt and select the one that is the most meaningful. These pairwise preferences are used to compute an Elo score (see Sec. B).
@@
Reference $20.2\pm 0.1$ $25.2\pm 0.1$ $4.02\pm 0.11$ $2180\pm 30$
RQ-transformer 8 RVQ 1.0 $\times 1.0$ $\times 1.0$ 26.7% $52.4\pm 0.2$ $36.3\pm 0.2$ $2.42\pm 0.12$ $1841\pm 25$ 4
RQ-transformer 8 RVQ 0.8 $\times 1.0$ $\times 1.0$ 26.7% $26.8\pm 0.1$ $33.1\pm 0.1$ $2.75\pm 0.14$ $1870\pm 30$ 3
CALM - Consistency - 1 step 1.0 $\times 1.3$ $\times 12.3$ 2.9% $42.9\pm 0.2$ $34.3\pm 0.2$ $2.82\pm 0.13$ $1947\pm 28$ 2
CALM - Consistency - 1 step 0.8 $\times 1.3$ $\times 12.3$ 2.9% $\bm{23.8}\pm\bm{0.1}$ $\bm{31.2}\pm\bm{0.1}$ $\bm{3.45}\pm\bm{0.14}$ $\bm{2023}\pm\bm{27}$ $\bm{1}$
@@ 5.2 Music Continuation
VAE: Our variational autoencoder (VAE) and codec architecture are adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz. We trained it to compress 32kHz mono music with a 25Hz frame rate. We experiment with bottleneck sizes of 96 and 128 dimensions. For comparison, MusicGen’s EnCodec model (Copet et al., 2023) also operates at 32kHz but uses a 4-level RVQ at 50Hz. In Tab. 3, we report reconstruction metrics (audio ViSQOL Chinen et al. (2020) and SISNR), showing that a 32-dim VAE matches MusicGen’s codec, and that at least 96 dimensions are needed for our VAE to match the quality of the 32-level RVQ configuration.
EnCodec Copet et al. (2023) 4 RVQ 50 2.2 2.41 5.62
VQ-VAE (inspired from Mimi) 32 RVQ 25 8.8 3.63 9.61
VAE 32 dims 25 2.23 5.51
VAE 96 dims 25 3.65 9.76
VAE 128 dims 25 4.01 10.3
@@
Model: Our music CALM builds on the MusicGen Medium backbone, a 1.35B parameter Transformer (see Sec. C for all the hyperparameters). We compare our method against: (1) Two discrete models using 32 and 16 RVQ codecs, each employing the same 1.35B backbone and a RQ-transformer for parallel codes prediction; (2) a MusicGen Medium variant trained with EnCodec without textual conditioning, using the same backbone and a delay pattern for codebook interleaving. Both MusicGen Medium and EnCodec models were trained on our dataset.
Results and Ablation: In Tab. 4, we report both objective metrics and the results of a human evaluation study for the task of music continuation, conditioned on a 3-second prompt. We compute the speed-up compared to the RQ-Transformer 32 RVQ, the Fréchet Audio Distance (FAD) which is the distance between Gaussian distributions fitted on VGG-obtained embeddings of the real and generated samples. We compute it on 4,000 model-generated continuations from the test set. The Acoustic Quality is a MOS score between 1 and 5. The Enjoyment metric is an Elo score (see Sec. B), computed by making human rater choose their favorite music out of generated pairs with the same 3s prompt. The evaluation protocol uses the same amount of annotators and examples as the speech study. We observe that CALM with consistency outperforms the 32 RVQ RQ-Transformer baseline on computed and human metrics while being $\times 1.9$ to $\times 2.2$ times faster for the overall speedup. While the RQ-Transformer takes $57.7\%$ of the inference time for the baseline, the consistency head only takes $6.6\%$ to $20.1\%$. As well, we train a CALM model with a TrigFlow head instead of consistency and it outperforms all the models but to the price of a slow inference.
@@
Scalability: We show in Sec. A.2 that CALM does improve with a bigger 3B parameters backbone. However, we leave a complete scalability study for future work.
Necessity of Consistency for a fast inference: We show in Tab. 6 that the consistency framework largely outperforms the TrigFlow framework for 10 inference steps and less (the regime where the Real time factor is greater than 1).
Table 6: Generation efficiency of TrigFlow CALM and Consistency CALM models. We compute the inference time to generate 30 seconds of audio, the corresponding Real Time Factor (RTF) as well as the FAD metric for different numbers of inference steps. For less than 10 steps (i.e. when streaming applications are possible, the consistency method outperforms the TrigFlow method).
# Steps Time (s) RTF FAD (TrigFlow) FAD (Consistency)
1 16.7 $\times 1.80$ $0.83\pm 0.04$
4 20.4 $\times 1.47$ $28.83\pm 0.20$ $0.71\pm 0.05$
10 27.7 $\times 1.08$ $4.62\pm 0.07$ $0.73\pm 0.05$
25 46.1 $\times 0.65$ $0.79\pm 0.04$ $0.96\pm 0.06$
50 76.8 $\times 0.39$ $0.74\pm 0.05$ $1.46\pm 0.06$
100 136.4 $\times 0.22$ $0.64\pm 0.04$ $2.05\pm 0.07$
5.3 Effect of gaussian temperature sampling
We evaluate how our proposed gaussian temperature sampling method affects acoustic diversity for consistency models, in comparison with temperature sampling on its discrete counterpart. More precisely, we compute WavLM speaker embeddings (Chen et al., 2021) over 100 unprompted speech generations with different values of temperature, both for an RQ-transformer and CALM. In both cases, the inner monologue text stream is generated with 0.8 temperature to ensure text diversity. We then average the pairwise cosine similarities of these embeddings, as a measure of diversity: the higher the average similarity, the lower speaker diversity in generated audio. The reference speaker similarity number is computed over 100 examples from the ground truth dataset. Fig. 3 shows, as expected, that speaker similarity tends to decrease with temperature, which means that temperature increases diversity with a similar trajectory to the discrete models.

### v3

@@ 3. Consistency‑Model Head
$\mathcal{L}_{\text{CALM}}(\theta,\phi,\psi)=\sum_{s=1}^{S}\mathbb{E}_{t,\bm{\epsilon}}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-F_{\bar{\phi}}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-\cos(t)\frac{df_{\bar{\phi}}(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s})}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right],$ (3)
@@ 4.3 Combining noisy long-term context and clean short-term context
In preliminary experiments, music generation models trained with the MAR framework were diverging quickly during inference because they were not robust to error accumulations. Applying noise injection during training slightly improves model stability but often reduces detail and instrument diversity, typically preserving only rhythmic elements, with audio fading into silence after 10–15 seconds. Since Pasini et al. (2024b) targets short, single-instrument clips, it’s unsurprising the method performs best in that constrained setting. We hypothesize that the added noise inhibits the backbone transformer from encoding fine-grained information into $\mathbf{z}^{s}_{\text{long}}$, limiting the MLP’s ability to reconstruct detailed audio. However, combining this with a short-context transformer computing $\mathbf{z}^{s}_{\text{short}}$ yields the best results (Tab. 5), likely because the clean short-term context restores local detail needed to model the distribution of the next $\mathbf{x}^{s}$.
@@ 4.5 Gaussian temperature sampling
Sampling strategies, such as temperature sampling, have a significant impact on generation quality in the discrete setup, particularly for speech. To replicate this behavior in the continuous domain, we introduce a sampling heuristic that results in comparable gains. Similarly to the GAN noise truncation trick presented in  Brock et al. (2018), we sample more from the high probability zone of the Gaussian to trade diversity for fidelity.
While the GAN truncation trick truncates the Gaussian noise such that values outside of a certain range are redrawn, we chose to reduce the variance of the Gaussian noise instead. This is mathematically equivalent to applying a temperature $\tau$ to the Gaussian if we change the standard deviation to $\sqrt{\tau}$. This makes temperature values between the discrete and continuous setups somewhat comparable, and we found that using a temperature of .8 for speech continuation was bringing good results in both setups. The effects of gaussian temperature are further discussed in Section C.
4.6 Latent Classifier Free Guidance
Classifier Free Guidance (CFG) (Ho and Salimans, 2022) is known to improve the generation quality of conditioned generative models. It can be applied for diffusion and flow matching models on the sampling trajectory as well as on the logits of autoregressive language models (Kreuk et al., 2023). Since CFG cannot be applied on the trajectory of 1-step consistency models we decide to apply the CFG on the outputs of the Backbone and Short-Context transformers. Formally, given $\mathbf{C}$ a conditioning and $\alpha$ the CFG coefficient, we compute for every $s$ of the sequence $\mathbf{Z}^{s}_{\text{CFG}}=\mathbf{Z}^{s}_{\emptyset}+\alpha(\mathbf{Z}^{s}_{C}-\mathbf{Z}^{s}_{\emptyset})$ and then we generate $\hat{\mathbf{x}}^{s}$ with the consistency head conditioned on $\mathbf{Z}^{s}_{\text{CFG}}$. We call this method Latent CFG, as it operates on the latent variable $\mathbf{Z}^{s}$ instead of the model output. It has been introduced in the video-to-audio model SoundReactor (Saito et al., 2025).
4.7 Latent Distillation
Once a teacher model has been trained and the desired classifier free guidance (CFG) coefficient has been selected for inference, we distill the CFG-guided teacher into a student model to avoid the need to double the batch size during inference when using CFG. To this end, we distill only the backbone transformer and directly copy the teacher’s MLP head into the student.
The distillation objective for the student backbone is an $\ell_{2}$ loss between the latent representation $\mathbf{Z}^{s}_{\text{distill}}$ produced by the student backbone and the CFG-guided latent representation $\mathbf{Z}^{s}_{\text{CFG}}$ produced by the teacher. Additionally, the student backbone transformer may contain fewer layers than the teacher. In practice for Pocket TTS, we distill a text-to-speech model with 24 transformer layers into a student model with only 6 layers, using a latent CFG coefficient of $\alpha=1.5$. See Sec. F for more details.
@@ 5.1 Speech Continuation
Table 1: Speech compression models. Our VAE is on par with the VQ-VAE on acoustic quality (MOSNeT) and outperforms it on semantic discriminability (ABX), PESQ Rix et al. (2001) and STOI Taal et al. (2011) and a MUSHRA for acoustic quality.
Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbit/s) MOSNet ($\uparrow$) ABX ($\downarrow$) PESQ ($\uparrow$) STOI ($\uparrow$) Acoustic Qual. ($\uparrow$)
VQ-VAE (Mimi) 8 RVQ 12.5 1.1kbps 3.11 9.4% 2.13 0.87 57.7 $\pm$ 1.3
VAE 32 dims 12.5 - 3.15 8.1% 2.42 0.90 66.0 $\pm$ 1.4
@@
Reference $20.2$ $25.2$ $4.02\pm 0.11$ $2180\pm 30$
RQ-transformer 8 RVQ 1.0 $\times 1.0$ $\times 1.0$ 26.7% $52.4$ $36.3$ $2.42\pm 0.12$ $1841\pm 25$ 4
RQ-transformer 8 RVQ 0.8 $\times 1.0$ $\times 1.0$ 26.7% $26.8$ $33.1$ $2.75\pm 0.14$ $1870\pm 30$ 3
CALM - Consistency - 1 step 1.0 $\times 1.3$ $\times 12.3$ 2.9% $42.9$ $34.3$ $2.82\pm 0.13$ $1947\pm 28$ 2
CALM - Consistency - 1 step 0.8 $\times 1.3$ $\times 12.3$ 2.9% $\bm{23.8}$ $\bm{31.2}$ $\bm{3.45}\pm\bm{0.14}$ $\bm{2023}\pm\bm{27}$ $\bm{1}$
Results: Tab. 1 shows that our 32-dimensional VAE matches an 8-RVQ Mimi codec on MOSNet (Lo et al., 2019), which measures audio quality, and exceeds it on the ABX metric (Schatz et al., 2013). ABX evaluates phonetic discriminability by testing whether a word like “bat” is represented closer to another “bat” utterance than to a similar-sounding word like “bit”, based on latent distances. Tab. 2 shows that the 1-step Consistency model outperforms the RQ-Transformer with 8 RVQ on all our automatic and human based metrics as well as on speed. For automatic metrics, we compute PPX and VERT as introduced by Lakhotia et al. (2021). The PPX metric measures the semantic meaningfulness of the generated speech. To do so we generate 1000 excerpts of speech of 30 second, we use Whisper (Radford et al., 2022) to compute textual transcriptions and finally compute the negative log-likelihood of the text tokens with a Mistral 7B LLM Jiang et al. (2023) and convert it to Perplexity. Because a model that generates poorly diverse but good quality sentences would perform well on the PPX metric, the authors of (Lakhotia et al., 2021) introduce the VERT metric (for diVERsiTy) which is a geometric mean of self- and auto-BLEU metrics. We use the official implementation from the fairseq Ott et al. (2019) repository.
To assess perceptual quality, we conduct two human evaluation studies involving 50 participants and 50 randomly selected examples from the English test set. Each participant rates 10 examples across the following evaluation protocols: For Acoustic Quality, participants are presented with all model continuations for the same prompt, including the ground truth reference, and rate the acoustic quality of each continuation on a 1 to 5 scale. For Meaningfulness, participants are shown two continuations of the same prompt and select the one that is the most meaningful. These pairwise preferences are used to compute an Elo score (see Sec. G).
Temperature Sampling and speaker similarity: In Sec. C, we show that our gaussian temperature sampling heuristic has similar effects on speaker similarity than the temperature sampling of the discrete model.
5.2 Text-to-Speech (TTS)
VAE, model and dataset: We use the same VAE as for speech continuation. Our TTS CALM builds on a 300M backbone transformer and uses the same architecture for the 10M parameters consistency sampling head. The text is fed to the backbone as a prefix with SentencePiece model (Kudo and Richardson, 2018) with a vocabulary size of 4k. The training data is a mix of public datasets totalizing to 88k hours of speech that is detailed in Sec. D.
Results: We evaluate on the Librispeech test-clean set using the same protocol as F5-TTS (Chen et al., 2025). We compare against four baselines: F5-TTS (Chen et al., 2025), DSM (Zeghidour et al., 2025), DiTAR (Jia et al., 2025), and SALAD (Turetzky et al., 2024). For DiTAR and SALAD, we report the paper results since the models are closed-source. We report Word Error Rate (WER) and Character Error Rate (CER) using Whisper-large-v3 (Radford et al., 2022), Speaker Similarity using WavLM-large (Chen et al., 2021b) as well as the results of a MUSHRA test for acoustic quality and a pairwise audio test for speaker similarity.
Table 3: Text-to-Speech models. Our CALM model with 1-step LSD outperforms baselines on WER, CER and Acoustic Quality. Results for Pocket TTS are in Sec. F.
Model Num. parameters WER CER SIM Acoustic Speaker SIM
($\downarrow$) ($\downarrow$) ($\uparrow$) Quality ($\uparrow$) Human ELO ($\uparrow$)
Reference 2.23 0.69 61.8 $\pm$ 2.4 1953 $\pm$ 24
Reference (with VAE) 0.57
F5 TTS (NFE=32) (Chen et al., 2025) 336M 2.42 0.66 54.7 $\pm$ 2.8 2032 $\pm$ 18
DSM (16 RVQ CFG=3 750M 1.95 0.67 60.2 $\pm$ 2.4 2112 $\pm$ 20
w.r.t text and audio prompt) (Zeghidour et al., 2025)
DiTAR (NFE=10) (Jia et al., 2025) 600M 2.39 0.67
SALAD (NFE=20) (Turetzky et al., 2024) 350M 0.74 0.54
CALM w/ LSD (NFE=1, CFG=1.5 w.r.t text) 313M 1.81 0.57 0.52 61.1 $\pm$ 2.3 1966 $\pm$ 23
Our CALM model with 1-step LSD (Boffi et al., 2025) outperforms baselines on WER, CER and Acoustic Quality but obtains a low speaker similarity score. This can be partially explained by the fact that when computing the speaker similarity between the reference prompt and the reference utterance that goes through the VAE (the second line of the tab) we obtain a similarity of 0.57. Yet, we demonstrate in Tab. 1 that our VAE faithfully reconstructs speech. Due to this surprising behavior, we decide to measure speaker similarity with a human study. We observe that all measured methods beat the ground truth, which means that they preserve well the voice of the audio prompt.
@@ 5.3 Music Continuation
VAE: Our variational autoencoder (VAE) and codec architecture is adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz. We trained it to compress 32kHz mono music with a 25Hz frame rate. Details and metrics are described in Sec. B.
@@
Model: Our music CALM builds on the MusicGen Medium backbone, a 1.35B parameter Transformer (see Sec. H for all the hyperparameters). We compare our method against: (1) Two discrete models using 32 and 16 RVQ codecs, each employing the same 1.35B backbone and a RQ-transformer for parallel codes prediction; (2) a MusicGen Medium variant trained with EnCodec without textual conditioning, using the same backbone and a delay pattern for codebook interleaving. Both MusicGen Medium and EnCodec models were trained on our dataset.
Results and Ablation: In Tab. 4, we report both objective metrics and the results of a human evaluation study for the task of music continuation, conditioned on a 3-second prompt. We compute the speed-up compared to the RQ-Transformer 32 RVQ, the Fréchet Audio Distance (FAD) which is the distance between Gaussian distributions fitted on VGG-obtained embeddings of the real and generated samples. We compute it on 4,000 model-generated continuations from the test set. The Acoustic Quality is a MOS score between 1 and 5. The Enjoyment metric is an Elo score (see Sec. G), computed by making human rater choose their favorite music out of generated pairs with the same 3s prompt. We observe that CALM with consistency outperforms the 32 RVQ RQ-Transformer baseline on computed and human metrics while being $\times 1.9$ to $\times 2.2$ times faster for the overall speedup. While the RQ-Transformer takes $57.7\%$ of the inference time for the baseline, the consistency head only takes $6.6\%$ to $20.1\%$. As well, we train a CALM model with a TrigFlow head instead of consistency and it outperforms all the models but to the price of a slow inference.
@@
Necessity of Consistency for a fast inference: We show in Sec. E.3 that the consistency framework largely outperforms the TrigFlow framework for 10 inference steps and less (the regime where the Real time factor is smaller than 1).
Scalability: We show in Sec. E.4 that CALM does improve with a bigger 3B parameters backbone. However, we leave a complete scalability study for future work.
5.4 Text-to-music generation
We compare CALM on the task of text-to-music generation. Since our dataset do not have any text labeling, we use CLAP (Elizalde et al., 2023) as a prefix conditioning and train MusicGen, our RQ-Transformer 32 RVQ baseline as well as CALM. During training, we drop the conditioning $20\%$ of the time in order to perform Classifier Free Guidance (Ho and Salimans, 2022) at inference. We use CLAP in audio mode when training. We use the same training data, VAE/codecs and hyperparameters as the models reported in Tab. 4. In Tab. 6, we show the results of our CLAP conditioned model evaluated on our internal test set with the CLAP conditioning being used in audio mode. As well, we report the results of our model and several baselines on the text-to-music benchmark on the MusicCaps dataset Agostinelli et al. (2023) in Sec.E.6.
Table 6: Text-to-music generation models on our test set where CLAP is used in audio mode.
Model Overall Sampler % time in FAD ($\downarrow$) Acoustic Enjoyment
Reference 3.75 $\pm$ 0.15 2104 $\pm$ 24
RQ-transformer 32 RVQ (CFG=3) $\times$ 1.0 $\times$ 1.0 57.7% 0.93 $\pm$ 0.07 3.16 $\pm$ 0.15 1998 $\pm$ 24 1
CALM - Consistency - 4 Steps (CFG=2) $\times$ 1.9 $\times$ 5.4 20.1% 0.91 $\pm$ 0.08 3.11 $\pm$ 0.14 1998 $\pm$ 24 1
MusicGen Medium $\times$ 1.3 0.0% 1.93 $\pm$ 0.14 2.54 $\pm$ 0.14 1946 $\pm$ 25 3